Question 1 3 pts Point charges of +6.83 μC and-1.26 μC are placed on the X-axis, at X 8.00 m and X 16.0 m, respectively. What charge must be placed at x 24.0 m, so that any charge placed at the origin would experience no electrostatic force? Specify your answer in uC, up to 1 decimal place.
Answer #1

let a unit charge be placed at origin .

let q  be the required charge to placed at x = 24

net force on the unit charge due to all three charges must be zero.

so,

k*6.83*10-6*1 / (8)2 + k*(-1.26*10-6) /(16)2 + k*q*1/(24)2 = 0

0.1067*10-6 - 0.0049*10-6 + q/(24)2 = 0

q = -58.63*10-6 C

q = -58.63 \mu C